OpenStudy (anonymous):

Which equations are represented by the graph? Choose all answers that are correct.

OpenStudy (anonymous):

On the graph the line is on (2,-4) 4x – 2y = 8 y – 4 = 2x y = 2x – 4 y = 1/2x - 4

OpenStudy (anonymous):

please help!

OpenStudy (anonymous):

Well first lets get them all in slope-intercept form

OpenStudy (anonymous):

okay

OpenStudy (anonymous):

Is there any way you can show the graph?

OpenStudy (anonymous):

give me a second

OpenStudy (anonymous):

Ok

OpenStudy (anonymous):

The slope goes up 2 and over 1, so we can go ahead and eliminate the last choice

OpenStudy (anonymous):

alright so is it a and c?

OpenStudy (anonymous):

That's correct

OpenStudy (anonymous):

AWESOME!! thank you so much

OpenStudy (anonymous):

You're welcome!